The area of a rectangle is 28 square meters. If the length is 2 times the width, then find the length and width of the rectangle. Round off your answers to the nearest hundredth
Answer by stanbon(75887)

The area of a rectangle is 28 square meters. If the length is 2 times the width, then find the length and width of the rectangle. Round off your answers to the nearest hundredth
--------
Let width = x
Then length = 2x
-----
Equation:
x*(2x) = 28 m^2
2x^2 = 28
x^2 = 14
x = sqrt(14)m that is the width
2x = 2*sqrt(14)m that is the length
------------
Cheers,
Stan H.
-------------------------